A parabola is tangent to the x-axis, has a y-intercept of 18 and passes through the point (2. 8). What are the two possible quadratic equations?
Answer by Alan3354(69443)

A parabola is tangent to the x-axis, has a y-intercept of 18 and passes through the point (2. 8). What are the two possible quadratic equations?
-------------------------
Since it's tangent, the 2 roots are the same.
(x - a)^2 = 0
x^2 - 2ax + a^2 = 0
a^2 = 18
a = +3sqrt(2)
a = -3sqrt(2)
--------------
--> y = x^2 ± 6sqrt(2) + 18
--------------------
One goes thru (2,8), the other goes thru (-2,8).